The specific heat capacity of silver is 0.24 J/°C·g.

Calculate the energy required to raise the temperature of 154.0 g Ag from 274 K to 298 K.
Energy =
J

Calculate the energy required to raise the temperature of 1.0 mole of Ag by 1.0°C (called the molar heat capacity of silver).
Molar Heat Capacity =
J/°C·mol

It takes 1.27 kJ of energy to heat a sample of pure silver from 11.3°C to 14.8°C. Calculate the mass of the sample of silver.
Mass =
g
1
Expert's answer
2020-06-08T15:30:54-0400

Q = cmΔT

Q = 0.24 * 154.0 * (298 - 274) = 887.04 J


M (Ag) = 107.87 g/mol

C = 0.24 J/°C·g - for 1 g of Ag

For 107.87 g (1 mol):

0.24 * 107.87 = 25.89 J/°C·mol


m = Q / cΔT

ΔT = 14.8 - 11.3 = 3.5

m = "1.27*1000 \/ (0.24 * 3.5)"

m = 1511.9 g